# Inverted-circuit zero-noise extrapolation for quantum gate error mitigation

### v1

@@ III.2 Measuring the error strength
$=$ ${\rm tr}\bigl{\{}|0\rangle\langle 0|\leavevmode\nobreak\ ({\mathcal{E}}_{U^{\dagger}}{\mathcal{E}}_{U})\left(|0\rangle\langle 0|\right)\}$
$=$ ${\rm tr}\bigl{\{}{\mathcal{E}}^{\dagger}_{U^{\dagger}}\left(|0\rangle\langle 0|\right)\leavevmode\nobreak\ {\mathcal{E}}_{U}\left(|0\rangle\langle 0|\right)\bigr{\}}$
@@
Here, ${\mathcal{E}}_{U^{\dagger}}^{\dagger}$ denotes the adjoint quantum channel, generally defined as ${\rm tr}\left[A\leavevmode\nobreak\ {\mathcal{E}}(B)\right]={\rm tr}\left[{\mathcal{E}}^{\dagger}(A)B\right]$. In case of a depolarising error channel (which is self-adjoint and commutes with unitary channels), we obtain $\tilde{\rho}=\rho$, but, in general, the operator $\tilde{\rho}$ differs from $\rho$. However, we assume that it exhibits the same fidelity as $\rho$, see Appendix A for justifying this assumption. Then, we can decompose $\tilde{\rho}$ in the same way as $\rho$:
@@
where $\langle\psi|\tilde{\sigma}|\psi\rangle=0$. Using the decomposition of $\rho$ and $\tilde{\rho}$ in Eqs. (15,22) together with $\langle\psi|\sigma|\psi\rangle=\langle\psi|\tilde{\sigma}|\psi\rangle=0$, we arrive at:
@@
Figure 3: Noise strength $\epsilon_{a}$, see Eq. (24), as a function of the probability $P_{0}$ to measure every qubit in state $|0\rangle$ for different values of $a$, see Eq. (23). In low-noise regime, the choice of $a$ has no major effect on the value of the error strength.
@@
First, we note that the quantity $a$ is unknown, but expected to be small for incoherent errors, where the purity of $\sigma$ and $\tilde{\sigma}$ is low. Indeed, if the gate errors are described by Pauli channels (which, for CNOT gates, can be enforced by randomised compiling, see Sec. II.3), then $\sigma$ and $\tilde{\sigma}$ arise from a mixture of a large number of states originating from all possible errors occurring at each gate. Thus, $a={\rm tr}(\sigma\tilde{\sigma})\leq\frac{1}{2}[{\rm tr}(\sigma^{2})+{\rm tr}(\tilde{\sigma}^{2})]\ll 1$. In case of a depolarising error, e.g., we obtain $a=\frac{1}{2^{q}-1}$ where $q$ denotes the number of qubits.
@@
independent of $a$. Quite intuitively, the factor two in the denominator can be explained by the fact that the error of ${\mathcal{E}}_{U}$ is multiplied by a factor two after adding ${\mathcal{E}}_{U\dagger}$. Note that Eq. (25) is not true for coherent errors, where the assumption $a\ll 1$ is not fulfilled. (In this case, we have checked that $a\propto\frac{1}{\epsilon}$ for $\epsilon\to 0$.)
Coming back to the incoherent case, we take into account deviations from the asymptotic behaviour, Eq. (25), by assuming a reasonable value of $a$, e.g. $a=\frac{1}{2^{q}}$ (slightly larger than the result obtained for depolarising error, see above). If $P_{0}<\frac{1}{2^{q}}$, we set $a=P_{0}$ to avoid the cases of no solution or two solutions mentioned above. In total, we finally obtain the following estimate of the noise strength $\epsilon$ from the measured probability $P_{0}$ of detecting all $q$ qubits in state $|0\rangle$:
@@ III.3 Structure of IC-ZNE application
The method IC-ZNE consists of the following steps: first, we take the same circuits which we also use in standard ZNE, i.e., with noise scaling factors $\lambda=1,3$ and $5$ obtained from repeated CNOT gates, and measure the expectation value $\langle A\rangle$ for each of them. Second, the noise strength $\epsilon$ of each circuit is measured by adding its inverse as explained above. Finally, we plot $\langle A\rangle$ as a function of $\epsilon$ and extrapolate to $\epsilon\to 0$ using a linear fit. More details and examples will be shown in Sec. IV.2. It is noteworthy, that any noise scaling factor can be chosen for IC-ZNE in principle.
@@ IV.1 Sample Circuits
In this section, we will test standard ZNE and IC-ZNE with two different circuits on simulators and on quantum hardware. The first circuit is an example of Grover’s quantum search algorithm [35] with three qubits and 10 CNOT gates (after specifically tailoring it to conform to IBM’s quantum device architecture, here ibmq_ehningen, through transpilation, see Appendix B. We consider the observable
@@
corresponding to the probability to measure one of the two solutions 101 or 011 of the search problem encoded by the oracle of our circuit. Without errors, we therefore obtain $\langle A_{\rm Grover}\rangle_{\rm ideal}=1$.
The second circuit is taken from Harrow-Hassidim-Lloyd algorithm (HHL algorithm) [36]. This algorithm can be used to solve linear systems of equations defined by a matrix $B\in\mathbb{C}^{N\times N}$ and a vector $\vec{b}\in\mathbb{C}^{N}$ to find $\vec{x}\in\mathbb{C}^{N}$, so that $B\vec{x}=\vec{b}$. We use an example with $N=2$ as described in [37] and
@@ IV.2 Fitting procedure and noise scaling
[…]
@@
Figure 4: Standard ZNE (a) and Inverted-Circuit ZNE (b) for the Grover circuit performed on IBM’s quantum device ibmq_ehningen. For each noise scaling factor $\lambda=1$ (green diamonds), $\lambda=3$ (blue triangles) and $\lambda=5$ (grey pentagons), we execute 16 different randomly twirled quantum circuits, each with 625 shots (amounting to 10 000 shots per scaling factor). (a) In standard ZNE, the measured expectation values $\langle A_{\rm Grover}\rangle$ are plotted against the noise scaling factor $\lambda$. A first order polynomial fit yields the extrapolated value $\langle A_{\rm Grover}\rangle=0.94$, which is significantly smaller than the ideal value $\langle A_{\rm Grover}\rangle_{\rm ideal}=1$ (horizontal dotted line). (b) In Inverted-Circuit ZNE, the error strength $\epsilon$ of each circuit is measured using the method of inverted circuits (with 625 additional shots per circuit). The same expectation values $\langle A_{\rm Grover}\rangle$ as in (a) are now plotted as a function of $\epsilon$ instead of $\lambda$, yielding a more accurate extrapolation result of $\langle A_{\rm Grover}\rangle=0.99$.
@@
Figure 5: Same as Fig. 4, but for the HHL circuit instead of the Grover circuit. Again, the extrapolation using Inverted-Circuit ZNE (b) yields better agreement of the extrapolated value $\langle A_{\text{HHL}}\rangle$ ($=0.57$ for sZNE and $0.63$ for IC-ZNE) with the exact value ($5/8=0.625$) than standard ZNE (a). Note that, due to a larger number of CNOT gates, the measured values of the error strengths $\epsilon$ are larger than in case of the Grover circuit, see Fig. 4b).
@@
In Figs. 4 and 5, the above procedure is illustrated for the case of the Grover and the HHL circuit, respectively, executed on IBM’s device ibmq_ehningen. The corresponding transpiled circuits are displayed in Appendix B, whereas the average error rates of the individual CNOT gates, as obtained from the calibration data given by IBM, are shown in Appendix C. From Figs. 4 and 5, we see that, in both cases, the IC-ZNE yields a more accurate extrapolation than the standard ZNE. This can be explained by the fact that the error strengths of the scaled circuits deviate from the simple scaling behaviour assumed in standard ZNE (see Sec. II.2). To illustrate this point, we plot in Fig. 6, for the same data as shown in Fig. 4 (Grover circuit), the scaling of the error strengths determined by the method of inverted circuits as a function of the noise scaling factor $\lambda$. Remember that the factor $\lambda$ defines the scaled circuits by replacing each CNOT gate with $\lambda$ CNOT gates. Standard ZNE assumes that the error of the circuit then increases by the same factor. In contrast, the error strengths, as measured by the method of inverted circuits, clearly deviate from this simple linear scaling. This demonstrates the benefits of IC-ZNE, which replaces the linear scaling assumption with measured error strength values.
@@
Figure 6: Scaled error strength $\epsilon/\epsilon_{0}$ (mean values and standard deviation obtained from 16 randomly twirled circuits) as a function of the noise scaling factor $\lambda$ for the same data as in Fig. 4 (Grover). The scaled error strength is normalised such that the mean value $\langle\epsilon\rangle/\epsilon_{0}=1$ for $\lambda=1$. A clear deviation from linear scaling ($\epsilon/\epsilon_{0}=\lambda$, dashed line) assumed in standard ZNE is observed.
IV.3 Simulations with depolarising noise
@@
So far, we have presented the results of a single extrapolation run for both, the Grover and HHL circuits. To confirm the robustness of our conclusions, we will now present the results of a statistical analysis obtained from repeating the above procedure 50 times. We will consider, both, runs performed on real quantum hardware (see Sec. IV.4 below) as well as simulations with depolarising noise of varying strength.
@@
Regarding the latter, we simulate the CNOT gate using the depolarising noise model. For simplicity, we assume that single-qubit gates, state preparation, and readout are error-free. This allows us to isolate the main source of error, which originates from CNOT gates in IBM’s superconducting devices. We examine the effect of increasing noise strength by choosing values of 1%, 2% and 5% average error rate per CNOT gate. Additionally, to analyse cases where error rates are unevenly distributed among different CNOT gates, we perform a simulation where all but one CNOT gate have low error rates (1%), whereas the remaining CNOT gate displays a particularly high error rate of 8%. After averaging over all CNOT gates executed in the respective circuit, this corresponds, in total, to 4% (Grover) or 4.9% (HHL) error per CNOT gate.
@@
The results of $M=50$ simulation runs per error model are shown in Fig. 7 (Grover) and Fig. 8 (HHL). In both cases, we perform sZNE and IC-ZNE. For comparison, also the raw data obtained from the unscaled circuit without performing ZNE is displayed. Note that randomised compiling is not employed in the simulations, since it has no impact on a depolarising channel. In each case, we obtain $M=50$ extrapolated values $\langle A_{\rm Grover}\rangle_{i}$ and $\langle A_{\text{HHL}}\rangle_{i}$ from the different simulation runs ($i=1,2,\dots,M)$. From these values, we generate a box plot showing the median value (central line) and the interquartile range (box). The whiskers extend to the highest and lowest values within 1.5 times the interquartile range. Any data points outside of these boundaries are considered outliers and are plotted separately as circles. The accompanying plots at the bottom display the root mean square deviation from the exact value:
@@
Figure 7: Comparison of inverted-circuit ZNE (IC-ZNE), standard ZNE (sZNE) and unmitigated raw data (Raw) for 50 simulation runs of the Grover circuit using a depolarising error model with different average error rates per CNOT gate: (a) 1%, (b) 2%, (c) 4% and (d) 5%. Whereas the errors are uniformly distributed across all CNOT gates in (a,b) and (d), (c) refers to an uneven error distribution, where only one CNOT gate displays a high error rate. For each method (Raw, sZNE and IC-ZNE), the obtained extrapolated values of $\langle A_{\rm Grover}\rangle$ are summarised in a box plot (see main text). The exact value $\langle A_{\rm Grover}\rangle_{\rm ideal}$ is shown by a dashed horizontal line. The root mean square error (RMSE), indicating the deviation from the exact value, is shown in the accompanying plot below on a logarithmic scale. The inverted-circuit ZNE (IC-ZNE) clearly performs better than the standard ZNE (sZNE), since it shows the smallest RMSE for all error rates.
Figure 8: Same as Fig. 7, but for the HHL instead of the Grover circuit. Again, our method (IC-ZNE) performs better than standard ZNE (sZNE). However, even for IC-ZNE, the obtained extrapolated values significantly differ from the exact value (horizontal line) in case of the largest two error rates (c) and (d). This indicates that, due to the larger number of CNOT gates as compared to the Grover circuit, the errors in (c) and (d) are too large even for IC-ZNE to work effectively.
From Figs. 7 and 8, it is evident that, both for the Grover and the HHL circuit, and for all four error models introduced above (with different error rates of the CNOT gates, evenly or unevenly distributed), IC-ZNE consistently outperforms sZNE, since it displays the smallest RMSE in all cases.
For the Grover circuit (Fig. 7), we observe that the values obtained with sZNE approach the raw values as the error rate increases from (a) to (d), resulting in the mitigation effect approaching zero. In contrast, the IC-ZNE remains effective also at larger error strengths, although a slight bias towards larger values of $\langle A_{\rm Grover}\rangle$ can be observed for the two highest error rates (c) and (d). This indicates a fundamental limitation of ZNE techniques, which cannot be expected to yield accurate results if the level of noise in the original circuit is too high.
This limitation becomes even more evident in case of the HHL circuit, see Fig. 7, which exhibits a larger number of CNOT gates than the Grover circuit (18 instead of 10, see Appendix B). We observe that at higher error rates, sZNE yields no improvement over raw values whatsoever, see Fig- 8(c) and (d). In contrast, IC-ZNE performs better, despite the extrapolated values deviating significantly from the exact values.
In cases where error rates differ among different CNOT gates, the results for Grover and HHL are comparable to those of circuits with a uniform distribution of the same (or similar) total error, see Fig. 7(c,d) and Fig. 8(c,d).

### v3

@@ III.2 Measuring the error strength
$=$ ${\rm tr}\bigl{\{}|0\rangle\langle 0|~{}({\mathcal{E}}_{U^{\dagger}}{\mathcal{E}}_{U})\left(|0\rangle\langle 0|\right)\}$
$=$ ${\rm tr}\bigl{\{}{\mathcal{E}}^{\dagger}_{U^{\dagger}}\left(|0\rangle\langle 0|\right)~{}{\mathcal{E}}_{U}\left(|0\rangle\langle 0|\right)\bigr{\}}$
@@
Here, ${\mathcal{E}}_{U^{\dagger}}^{\dagger}$ denotes the adjoint quantum channel, generally defined as ${\rm tr}\left[A~{}{\mathcal{E}}(B)\right]={\rm tr}\left[{\mathcal{E}}^{\dagger}(A)B\right]$. The state $\tilde{\rho}$ is called ”dual state” in Refs. [21, 22], and the same circuit that we propose in order to measure the error strength has been used there to determine the purity ${\rm Tr}(\rho\tilde{\rho})$, see Eq. (22), needed to measure the purified expectation value $\langle A\rangle={\rm Tr}\{A\rho\tilde{\rho}\}/{\rm Tr}\{\rho\tilde{\rho}\}$.
In case of a depolarizing error channel (which is self-adjoint and commutes with unitary channels), we obtain $\tilde{\rho}=\rho$, but, in general, the operator $\tilde{\rho}$ differs from $\rho$. However, we assume that it exhibits the same fidelity as $\rho$, see Appendix A for justifying this assumption. Then, we can decompose $\tilde{\rho}$ in the same way as $\rho$:
@@
where $\langle\psi|\tilde{\sigma}|\psi\rangle=0$. Using the decomposition of $\rho$ and $\tilde{\rho}$ in Eqs. (17, 24) together with $\langle\psi|\sigma|\psi\rangle=\langle\psi|\tilde{\sigma}|\psi\rangle=0$, we arrive at:
@@
Figure 3: Noise strength $\epsilon_{a}$, see Eq. (26), as a function of the probability $P_{0}$ to measure every qubit in state $|0\rangle$ for different values of $a$, see Eq. (25). In a low-noise regime, the choice of $a$ has no major effect on the value of the error strength.
@@
First, we note that the quantity $a$ is unknown, but expected to be small for incoherent errors, where the purity of $\sigma$ and $\tilde{\sigma}$ is low. Indeed, if the gate errors are described by Pauli channels (which, for CNOT gates, can be enforced by randomized compiling, see Sec. II.3), then $\sigma$ and $\tilde{\sigma}$ arise from a mixture of a large number of states originating from all possible errors occurring at each gate. Thus, $a={\rm tr}(\sigma\tilde{\sigma})\leq\frac{1}{2}[{\rm tr}(\sigma^{2})+{\rm tr}(\tilde{\sigma}^{2})]\ll 1$. In case of a depolarizing error, e.g., we obtain $a=\frac{1}{2^{q}-1}$ where $q$ denotes the number of qubits.
@@
independent of $a$. Quite intuitively, the factor two in the denominator can be explained by the fact that the error of ${\mathcal{E}}_{U}$ is multiplied by a factor of two after adding ${\mathcal{E}}_{U\dagger}$. Note that Eq. (27) is not true for coherent errors, where the assumption $a\ll 1$ is not fulfilled. (In this case, we have checked that $a\propto\frac{1}{\epsilon}$ for $\epsilon\to 0$.)
Coming back to the incoherent case, we take into account deviations from the asymptotic behavior, Eq. (27), by assuming a reasonable value of $a$, e.g. $a=\frac{1}{2^{q}}$ (slightly larger than the result obtained for depolarizing error, see above). If $P_{0}<\frac{1}{2^{q}}$, we set $a=P_{0}$ to avoid the cases of no solution or two solutions mentioned above. In total, we finally obtain the following estimate of the noise strength $\epsilon$ from the measured probability $P_{0}$ of detecting all $q$ qubits in state $|0\rangle$:
@@ III.3 Structure of IC-ZNE application
The method IC-ZNE consists of the following steps: first, we take the same circuits which we also use in standard ZNE, i.e., with noise scaling factors $\lambda=1,3$ and $5$ obtained from repeated CNOT gates, and measure the expectation value $\langle A\rangle$ for each of them. Second, the noise strength $\epsilon$ of each circuit is measured by adding its inverse as explained above. Finally, we plot $\langle A\rangle$ as a function of $\epsilon$ and extrapolate to $\epsilon\to 0$ using a linear fit. More details and examples will be shown in Sec. IV.2. It is noteworthy that any noise scaling factor can be chosen for IC-ZNE in principle.
@@ IV.1 Sample circuits
In this section, we will test standard ZNE and IC-ZNE with two different circuits on simulators and on quantum hardware. The first circuit is an example of Grover’s quantum search algorithm [43] with three qubits and 10 CNOT gates (after specifically tailoring it to conform to IBM’s quantum device architecture, here ibmq_ehningen, through transpilation, see Appendix B). We consider the observable
@@
corresponding to the probability of measuring one of the two solutions 101 or 011 of the search problem encoded by the oracle of our circuit. Without errors, we therefore obtain $\langle A_{\rm Grover}\rangle_{\rm ideal}=1$.
The second circuit is taken from the Harrow-Hassidim-Lloyd algorithm (HHL algorithm) [44]. This algorithm can be used to solve linear systems of equations defined by a matrix $B\in\mathbb{C}^{N\times N}$ and a vector $\vec{b}\in\mathbb{C}^{N}$ to find $\vec{x}\in\mathbb{C}^{N}$, so that $B\vec{x}=\vec{b}$. We use an example with $N=2$ as described in [45] and
@@ IV.2 Fitting procedure and noise scaling
In all cases (unless otherwise specified), we use the following parameters: the original circuits corresponding to $\lambda=1$ are the transpiled circuits shown in Appendix B. We generate two scaled circuits with scaling factors $\lambda=3$ and 5 by multiplying CNOT gates, as explained in Sec. II.2. If randomized compiling is employed, 16 different twirled versions of each circuit are randomly generated and executed with $10~{}000/16=625$ shots, from which the expectation values $\langle A\rangle$ are determined separately for each of the 16 twirled versions. In the case of IC-ZNE, the noise strength $\epsilon$ of each version is additionally determined, again using $625$ shots for every twirled circuit. The total number of shots thus amounts to 30 000 for sZNE (i.e., 10 000 shots per scaling factor $\lambda$) and 60 000 for IC-ZNE (due to the additional measurement of the error strength $\epsilon$). Thereby, we obtain in total $3\times 16=48$ data points $(\lambda,\langle A\rangle)$ (sZNE) or $(\epsilon,\langle A\rangle)$ (IC-ZNE), which are fitted by an exponential function or first-order polynomial to obtain the extrapolated noise-free value of $\langle A\rangle$ at $\lambda=0$ (sZNE) or $\epsilon=0$ (IC-ZNE), respectively.
@@
Figure 4: Standard ZNE (a) and Inverted-Circuit ZNE (b) for the Grover circuit performed on IBM’s quantum device ibmq_ehningen. For each noise scaling factor $\lambda=1$ (green diamonds), $\lambda=3$ (blue triangles) and $\lambda=5$ (grey pentagons), we execute 16 different randomly twirled quantum circuits, each with 625 shots (amounting to 10 000 shots per scaling factor). The observed variation in the data points for the different noise scaling factors can be attributed to the use of different twirling gates, as well as to statistical fluctuations and variations introduced by the employed backend. The error bars (standard deviation) of the extrapolated values at zero noise ($\lambda=0$ or $\epsilon=0$, respectively) are determined from the covariance matrix of the fit parameters returned by the curve fitting routine.(a) In standard ZNE, the measured expectation values $\langle A_{\rm Grover}\rangle$ are plotted against the noise scaling factor $\lambda$. An exponential fit yields the extrapolated value $\langle A_{\rm Grover}\rangle=0.9703\pm 0.02131$, which is significantly smaller than the ideal value $\langle A_{\rm Grover}\rangle_{\rm ideal}=1$ (horizontal dotted line). (b) In Inverted-Circuit ZNE, the error strength $\epsilon$ of each circuit is measured using the method of inverted circuits (with 625 additional shots per circuit). The same expectation values $\langle A_{\rm Grover}\rangle$ as in (a) are now plotted as a function of $\epsilon$ instead of $\lambda$, yielding a more accurate extrapolation result of $\langle A_{\rm Grover}\rangle=0.9943\pm 0.0085$ with a first order linear fit.
@@
Figure 5: Same as Fig. 4, but for the HHL circuit instead of the Grover circuit. Again, the extrapolation using Inverted-Circuit ZNE (b) yields better agreement of the extrapolated value $\langle A_{\text{HHL}}\rangle$ $=0.5940\pm 0.0252$ for sZNE and $0.6327\pm 0.0166$ for IC-ZNE) with the exact value ($5/8=0.625$) than standard ZNE (a). Note that, due to a larger number of CNOT gates, the measured values of the error strengths $\epsilon$ are larger than in case of the Grover circuit, see Fig. 4b).
@@
In Figs. 4 and 5, the above procedure is illustrated for the case of the Grover and the HHL circuit, respectively, executed on IBM’s device ibmq_ehningen. The corresponding transpiled circuits are displayed in Appendix B, whereas the average error rates of the individual CNOT gates, as obtained from the calibration data given by IBM, are shown in Appendix D. From Figs. 4 and 5, we see that, in both cases, IC-ZNE yields a more accurate extrapolation than standard ZNE: the extrapolated values at zero noise ($\lambda=0$ or $\epsilon=0$, respectively) are closer to the exact result, and their standard deviations (error bars) are smaller. This can be explained by the fact that the error strengths of the scaled circuits deviate from the simple scaling behavior assumed in standard ZNE (see Sec. II.2). To illustrate the deviation from the simple scaling behavior, we plot in Fig. 6, for the same data as shown in Fig. 4 (Grover circuit), the scaling of the error strengths $\epsilon/\epsilon_{0}$ determined by the method of inverted circuits as a function of the noise scaling factor $\lambda$, where $\epsilon_{0}$ denotes the original error strength of the unscaled circuit ($\lambda=1$). Remember that the factor $\lambda$ defines the scaled circuits by replacing each CNOT gate with $\lambda$ CNOT gates. Standard ZNE with exponential fit assumes the following behavior of $\langle A\rangle$ as a function of $\lambda$:
$\langle A\rangle=a_{1}e^{-a_{2}\lambda}+a_{3}$ (33)
(see also Sec. II.2), where the ideal value is reached at $\lambda=0$, i.e., $\langle A\rangle_{\rm ideal}=a_{1}+a_{3}$. Therefore, we can rewrite Eq. (33) as follows:
$\langle A\rangle=e^{-a_{2}\lambda}\langle A\rangle_{\rm ideal}+\left(1-e^{-a_{2}\lambda}\right)\langle A\rangle_{\rm noisy},$ (34)
where $\langle A\rangle_{\rm noisy}=a_{3}$ refers to the expectation value reached in the strong-noise limit ($\lambda\to\infty$). This equation can be interpreted as follows: the quantum state at noise scaling factor $\lambda$ is a mixture of the ideal state $\rho_{\rm ideal}=|\psi\rangle\langle\psi|$ and a noisy state $\rho_{\rm noisy}$ (reached in the limit of strong noise):
$\rho=e^{-a_{2}\lambda}|\psi\rangle\langle\psi|+\left(1-e^{-a_{2}\lambda}\right)\rho_{\rm noisy}.$ (35)
In particular, this is true for depolarizing noise, where $\rho_{\rm noisy}=\rho_{*}$, see Sec. II.2. Finally, Eq. (35) yields the following relation between the noise scaling factor $\lambda$ and the error strength $\epsilon$, see Eqs. (15,16):
$\epsilon=\Bigl{(}1-\langle\psi|\rho_{\rm noisy}|\psi\rangle\Bigl{)}~{}\Bigl{(}1-e^{-a_{2}\lambda}\Bigr{)}$ (36)
or
$\frac{\epsilon}{\epsilon_{0}}=\frac{1-e^{-a_{2}\lambda}}{1-e^{-a_{2}}},$ (37)
which we plotted in Fig. 6 (dashed line) as the scaling behaviour assumed in standard ZNE.
In contrast, the error strengths measured by the method of inverted circuits deviate significantly from this simple exponential scaling. This proves that the noise on the real quantum hardware cannot be adequately described by a depolarizing noise model (for which exponential scaling holds as shown above) and demonstrates the benefits of IC-ZNE, which replaces the exponential scaling assumption with measured error strength values.
@@
Figure 6: Scaled error strength $\epsilon/\epsilon_{0}$ (mean values and standard deviation obtained from 16 randomly twirled circuits) as a function of the noise scaling factor $\lambda$ for the same data as in Fig. 4 (Grover). The scaled error strength is normalized such that the mean value $\langle\epsilon\rangle/\epsilon_{0}=1$ for $\lambda=1$. A clear deviation from exponential scaling (dashed line), see Eq. (37) with $a_{2}=0.1350$ extracted from the exponential fit in Fig. 4(a), assumed in standard ZNE is observed.
IV.3 Simulations with depolarizing noise
@@
So far, we have presented the results of a single extrapolation run for both, the Grover and HHL circuits. To confirm the robustness of our conclusions, we will now present the results of a statistical analysis obtained from repeating the above procedure 50 times. We will consider, both: runs performed on real quantum hardware (see Sec. IV.4 below) as well as simulations with depolarizing noise of varying strength.
@@
Regarding the latter, the CNOT and single-qubit gates are simulated using the depolarizing noise model. For simplicity, we assume that state preparation, and readout are error-free. This allows us to isolate the main source of error, which originates from CNOT gates in IBM’s superconducting devices. We examine the effect of increasing noise strength by choosing values of 0.5%, 1%, 2% and 5% error rate per CNOT gate and 1/10 of the respective value for single qubit gates.
As discussed in Sec. IV.2, exponential scaling of the noise strength, as assumed in standard ZNE, can be proven to hold for fully depolarizing noise, where each noisy gate exhibits the same error channel given by Eq. (8) with constant $p$.
@@
The results of $M=50$ simulation runs per error model are shown in Fig. 7 (Grover) and Fig. 8 (HHL). In both cases, we perform sZNE and IC-ZNE. For comparison, the raw data obtained from the unscaled circuit without performing ZNE is also displayed. Note that randomized compiling is not employed in the simulations, since it has no impact on a depolarizing channel. In each case, we obtain $M=50$ extrapolated values $\langle A_{\rm Grover}\rangle_{i}$ and $\langle A_{\text{HHL}}\rangle_{i}$ from the different simulation runs ($i=1,2,\dots,M)$. From these values, we generate a box plot showing the median value (central line) and the interquartile range (box). The whiskers extend to the highest and lowest values within 1.5 times the interquartile range. Any data points outside of these boundaries are considered outliers and are plotted separately as circles. The accompanying plots at the bottom display the root mean square deviation from the exact value:
@@
Figure 7: Comparison of inverted-circuit ZNE (IC-ZNE), standard ZNE (sZNE) and unmitigated raw data (Raw) for 50 simulation runs of the Grover circuit using a depolarizing error model with different error rates per CNOT gate: (a) 0.5%, (b) 1%, (c) 2% and (d) 5%. The error rate for single qubit gates is one-tenth of the CNOT error rate. For each method (Raw, sZNE and IC-ZNE), the obtained extrapolated values of $\langle A_{\rm Grover}\rangle$ are summarized in a box plot (see main text). The exact value $\langle A_{\rm Grover}\rangle_{\rm ideal}$ is shown by a dashed horizontal line. The root mean square error (RMSE), indicating the deviation from the exact value, is shown in the accompanying plot below on a logarithmic scale. In all cases, the inverted-circuit ZNE (IC-ZNE) exhibits the smallest root mean square error.
Figure 8: Same as Fig. 7, but for the HHL instead of the Grover circuit. IC-ZNE and sZNE show similar RMSE values. For CNOT error rates $\leq$ 2%, the distribution for sZNE is broader than for IC-ZNE. IC-ZNE performs at a minimum of 1.6 times (in terms of RMSE) better than sZNE in (a), (b) and (c). For higher error rates, the obtained extrapolated values significantly differ from the exact value (horizontal line). This indicates that, due to the larger number of CNOT gates as compared to the Grover circuit, the errors in (d) are too large for ZNE to work effectively.
From Figs. 7 and 8, it is evident that, both for the Grover and the HHL circuit, and for all four error models introduced above (with different error rates of the CNOT and single-qubit gates), IC-ZNE demonstrates superior performance compared to sZNE, exhibiting the smallest RMSE in all but one instance (corresponding to HHL with largest error rate, see below).
For the Grover circuit (Fig. 7), we observe a slight bias of IC-ZNE towards larger values of $\langle A_{\rm Grover}\rangle$ for the two highest error rates (c) and (d). IC-ZNE performs at a minimum of 1.6 times better (in terms of RMSE) than sZNE. In the best case, see Fig. 7(b), IC-ZNE exhibits a 3.2-fold improvement over sZNE by comparison of the RMSE values.
For the HHL circuit, Figs. 8 (a), (b) and (c) demonstrate that sZNE exhibits a greater degree of variation than IC-ZNE. For the lowest error rate in (a), IC-ZNE performs 2.5 times better, because it has fewer variations and outliers. For CNOT error rates reaching 5%, see Fig. 8 (d), the extrapolated values of, both, sZNE and IC-ZNE deviate significantly from the exact value (horizontal line). This indicates a fundamental limitation of ZNE techniques, which cannot be expected to yield accurate results if the level of noise in the original circuit is too high (due to the larger number of CNOT gates in comparison to the Grover circuit).